OpenStudy (yinshy):

A bag contains a total of 25 marbles. Five of the marbles are blue, 3 are red, 8 are green, and 9 are white. If a marble is drawn from the bag and then replaced 50 times, about how many times would a green marble be drawn? A. 4 B. 8 C. 16 D. 24

OpenStudy (yinshy):

@calculusxy

OpenStudy (yinshy):

@jameshorton

OpenStudy (calculusxy):

I think we can make a proportion in here. Since it is without replacement, we won't have to worry about altering the denominators. From the given situation, there is a total of 25 marbles and 8 of them are green (which can be shown as \(\large \frac{total}{green} = \frac{25}{8}\). Similarly, for the 50 trials (total) we would like to figure out the probability of picking a green marble (unknown or x). That can be shown as \(\large \frac{50}{x}\). Now let's make a proportion from those two fractions: \(\large \frac{25}{8} = \frac{50}{x}\) Solve for \(x\).

OpenStudy (yinshy):

so c?

OpenStudy (calculusxy):

Correct!

OpenStudy (jameshorton):

sup